OpenStudy (anonymous):

-5x - 5 = -4x + 4. What is x?

OpenStudy (anonymous):

\[-5x-5=-4x+4\] right?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

takes exactly two steps a) add \(5x\) to both sides b) subtract \(4\) from both sides

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

No, sorry

OpenStudy (anonymous):

on the left you have \(-5x\) yes?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so lets get rid of it by adding \(5x\) to both sides it will be gone from the left side because \(-5x+5x=0\)

OpenStudy (anonymous):

So the right side has to equal 0 right

OpenStudy (anonymous):

So would it be 1?

OpenStudy (anonymous):

it would be \(x\) lets go slow

OpenStudy (anonymous):

\[-5x-5=-4x+4\] add \(x\) to both sides get \[-5x+5x-5=5x-4x+4\] that gives \[-5=x+4\] because \(5x-4x=x\)

OpenStudy (anonymous):

i meant "add \(5x\) to both sides"

OpenStudy (anonymous):

now that you have \[-5=x+4\] you and solve for \(x\) by subtracting \(4\) from both sides to get \(x\) by itself

OpenStudy (anonymous):

So would it be -9 = x ?

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

yw